# Compositional Coordination for Multi‑Robot Teams with Large Language Models

### v2

@@ II-B2 Team-Level Integration
Extending LLMs to multi-robot systems introduces challenges in coordination, communication, and scalability, with research still in the early stages exploring their use for flexible, robust multi-robot coordination [39]. Some works use LLMs to translate natural language into formal representations [40, 41, 42, 43], but they are limited to simple missions and cannot handle multi-robot teams or decomposable long-horizon tasks. Others propose prompt-based frameworks for multi-robot coordination using LLMs [27, 28]. For instance, Zhao et al. introduced RoCO [27], where robots coordinate via interactive dialogues. However, such frameworks face limitations with large robot teams (e.g., $\geq 10$) and complex language-based missions. Shyam et al. proposed SMART-LLM [28], an LLM-based task planner for multi-robot teams with task decomposition and allocation, but it lacks domain knowledge, limits generalization beyond in-context examples, and cannot handle motion-level language commands. Our work is closely related to GenSwarm [44], which translates natural language task descriptions into executable Python code. While both aim to enable language-driven multi-robot coordination, our framework differs in two key ways. First, our modular design supports a broader range of coordination problems, including multi-team tasks with interdependencies, unlike GenSwarm’s focus on simple, single-team scenarios. Second, we incorporate a structured knowledge base to guide the LLM, enhancing extensibility, whereas GenSwarm relies solely on in-context learning, limiting generality.
Figure 2: A comprehensive demonstration of LAN2CB.   user- or expert-provided content.   LLM-generated mission descriptions.   LLM-generated code.
@@ IV Method
    Mission Completion: Once the mission finish condition is met, the system terminates and the mission is formally marked as complete.
@@ IV-A Mission Analysis
node_type Parallel |||| Sequence |||| Action.
@@
status Idle |||| Running |||| Success |||| Failure.
@@
action_type visit_points |||| follow_targets |||| herd.
@@
Ready-to-Run Action Node(s) Selection: After constructing the complete behavior tree and integrating the dependency analysis results, we ask the LLM to identify action nodes that are immediately executable. Specifically, these are nodes for which all dependencies are satisfied and all preconditions are met in the current environment state. The ready-to-run action nodes are then selected and forwarded for downstream processing, code generation. This approach ensures that robot actions are initiated only when all prerequisite tasks and constraints have been resolved, thereby supporting safe and efficient mission execution. After selecting action nodes that are ready-to-run, we instruct the LLM to update the status of all nodes in the behavior tree, marking the selected nodes and their parent nodes as Running.
@@ V-A Performance Across Diverse Missions
Figure 4: Snapshots of real-world experiments.   idle,   running,   finished.
@@ V-B Impact of standardized template
To quantify the impact of using a template to standardize the mission descriptions on LAN2CB, we ablated the template-reformatting stage that aligns user-provided natural language mission descriptions to our predefined mission schema. As reported in Tab. II, this omission results in an average drop of $\sim 15\%$ in success rates and causes the error rate to become 2.14×2.14$\times$2.14 × more frequent across all mission categories. These results highlight the critical role of structured prompting in maintaining robust multi-robot performance.
V-C Real World Demonstrations

### v3

@@ II-B2 Team-Level Integration
Extending LLMs to multi-robot systems introduces challenges in coordination, communication, and scalability, with research still in the early stages exploring their use for flexible, robust multi-robot coordination [39]. Some works use LLMs to translate natural language into formal representations [40, 41, 42, 43], but they are limited to simple missions and cannot handle multi-robot teams or decomposable long-horizon tasks. Others propose prompt-based frameworks for multi-robot coordination using LLMs [27, 28]. For instance, Zhao et al. introduced RoCO [27], where robots coordinate via interactive dialogues. However, such frameworks face limitations with large robot teams (e.g., $\geq 10$) and complex language-based missions. Shyam et al. proposed SMART-LLM [28], an LLM-based task planner for multi-robot teams with task decomposition and allocation, but it lacks domain knowledge, limits generalization beyond in-context examples, and cannot handle motion-level language commands. Our work is closely related to GenSwarm [44], which translates natural language task descriptions into executable Python code. While both aim to enable language-driven multi-robot coordination, our framework differs in two key ways. First, our modular design supports a broader range of coordination problems, including multi-team tasks with interdependencies and trigger events, unlike GenSwarm’s focus on simple, single-team scenarios. Second, we incorporate a structured knowledge base to guide the LLM, enhancing extensibility, whereas GenSwarm relies solely on in-context learning, limiting generality.
Figure 2: A comprehensive demonstration of LAN2CB.   human-provided content.   LLM-generated mission descriptions.   LLM-generated code.
@@ IV Method
    Mission Completion: Once the mission finish condition is met, the mission is formally marked as complete.
@@ IV-A Mission Analysis
node_type Parallel $|$ Sequence $|$ Action.
@@
status Idle $|$ Running $|$ Success $|$ Failure.
@@
action_type visit_points $|$ follow_targets $|$ herd.
@@
Ready-to-Run Action Node(s) Selection: After constructing the complete behavior tree and integrating the dependency analysis results, we ask the LLM to identify action nodes that are immediately executable. Specifically, these are nodes for which all preconditions are met. The ready-to-run action nodes are then selected and forwarded for downstream processing, code generation. This approach ensures that robot actions are initiated only when all prerequisite tasks and constraints have been resolved, thereby supporting safe and efficient mission execution. After selecting action nodes that are ready-to-run, we instruct the LLM to update the status of all nodes in the behavior tree, marking the selected nodes and their parent nodes as Running.
@@ V-A Performance Across Diverse Missions
Figure 4: Realistic-setting experiments.   idle,   running,   finished.
@@ V-B Impact of standardized template
To quantify the impact of using a template to standardize the mission descriptions on LAN2CB, we ablated the template-reformatting stage that aligns user-provided natural language mission descriptions to our predefined mission schema. As reported in Tab. II, this omission results in an average drop of $\sim 15\%$ in success rates and causes the error rate to become $2.14\times$ more frequent across all mission categories. These results highlight the critical role of structured prompting in maintaining robust multi-robot performance.
Figure 5: Real-world experiments.   idle,   running,   finished.
V-C Realistic Simulated Scenarios
We validated our framework in three simulation scenarios designed to approximate real-world conditions (see Fig. 4).
V-D Real World Demonstrations